Find the equation for the parabola with focus at (3, -3) and vertex at (3, 1).
Answer by lwsshak3(11628)

Find the equation for the parabola with focus at (3, -3) and vertex at (3, 1).
***
given parabola opens downward.
Its basic equation: (x-h)^2=-4p(y-k), (h,k)=(x,y) coordinates of the vertex
Given vertex:(3,1)
axis of symmetry: x=3
p=4 (distance from vertex to focus on the axis of symmetry)
4p=16
Equation of given parabola:
(x-3)^2=-16(y-1)